OpenStudy (anonymous):

how do you find the derivate to the fraction -x^4/4

OpenStudy (kc_kennylau):

\[\frac d{dx}\left(-\frac{x^4}4\right)\]\[=-0.25\frac d{dx}x^4\] Hope you can continue :)

OpenStudy (anonymous):

so -x3 right?

OpenStudy (anonymous):

@khadishacicley

OpenStudy (kc_kennylau):

yes sorry for late reply

OpenStudy (anonymous):

-x^3 i mean

OpenStudy (kc_kennylau):

yes

OpenStudy (anonymous):

ok… i had a feeling i was doing it right but it keeps saying I'm wrong :/.. thank you for the help :)

OpenStudy (kc_kennylau):

no problem :)